OpenStudy (anonymous):

i need the equation of the line that has the indicated slope and contains the indicated point. final answer should be in standard form... m=4/3;(-4,-5)

OpenStudy (smarty):

y=mx+b

OpenStudy (smarty):

so it woul be y=4/3x+b, plug in your points to solve for "b" -5=-16/3+b -5+16/3=b 1/3=b

OpenStudy (anonymous):

i thought b was the y-value

OpenStudy (smarty):

The Y-value is separate...

OpenStudy (anonymous):

so how would that look in standard form? Ax+By=C

OpenStudy (smarty):

Oh lolol xD. I read the question wrong.

OpenStudy (anonymous):

\[y=mx+b,m=4/3\rightarrow y= (4/3)x+b\]Considering the point (-4,-5)\[-5=(4/3)\times(-4)+b \rightarrow b = 1/3\]Thus,\[y=(4/3)x+(1/3)\] Cheers!